### v2

@@ 1.1   Related literature
Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018)). We contribute to this literature in three important ways.
@@
The first contribution of this paper is to investigate the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH” in this section). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption. Hudson et al. (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings. Build on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs; the settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
@@
In this paper, we first formalize $2\times 2$ DID-IV designs and complement de Chaisemartin (2010) and Hudson et al. (2017). Specifically, we introduce the path of the instrument in potential outcomes and uncover the ”hidden assumption” in $2\times 2$ DID-IV designs; this requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument. We also investigate the selection mechanism behind $2\times 2$ DID-IV designs, and clarify the interpretation of the parallel trends assumption in the outcome.
@@
Given our identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs. The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicate the relative advantages of DID-IV over Fuzzy DID designs.
@@
The second contribution of this paper is to extend $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units. In practice, in order to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for the treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). No previous studies, however, extend $2\times 2$ DID-IV designs to such settings. Our staggered DID-IV designs ensure the theoretical validity of their identification strategies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
@@
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g. Black et al. (2005), Lundborg et al. (2014)). In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time. Our estimation method developed in this paper would serve as an alternative to the TWFEIV estimators and make DID-IV designs more credible in a given application.
@@
Our paper is also related to a recent and growing DID literature. Several recent papers have pointed out the issue of implementing DID designs via two-way fixed effects regressions or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al. (2021); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)). All of these papers have documented that regression coefficients by the conventional approaches may fail to properly summarize treatment effects under DID designs. Some of these papers have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity. None of these papers, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in the timing of policy adoption across units as an instrument for the treatment. In this paper, we formalize the underlying identification strategy as DID-IV designs, and provide the reliable estimation method. Built on the work in this paper, Miyaji (2024) studies the properties of the TWFEIV estimand under staggered DID-IV designs, and point out the issue of implementing this design via TWFEIV regressions.
@@ 2   DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in a two-period/two-group setting. We first establish the target parameter and identifying assumptions. We then uncover the selection mechanism behind the canonical DID-IV design, and clarify the interpretation of the parallel assumption in the outcome. We next describe the relationship between DID and DID-IV, and show that DID-IV is DID with the noncompliance of the assigned treatment. At the end of this section, we also describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
@@ 2.1   Set up
We introduce the notation we use throughout section 2. We consider a panel data setting with two periods and $N$ units. For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support. For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receives the treatment in period $t$. Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$. Throughout section 2, we assume that $\{Y_{i,0},Y_{i,1},D_{i,0},D_{i,1},Z_{i,0},Z_{i,1}\}_{i=1}^{N}$ are independent and identically distributed (i.i.d).
We introduce the path of the treatment and the instrument. Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ be the instrument path. We assume that no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$. Hereafter, we refer to this as a sharp assignment of the instrument. We denote $E_{i}\in\{0,1\}$ to be the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group). In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment: we assume that the treatment path can take four values with non-zero probability, that is, we have $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or a policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment. For instance, Duflo (2001) estimates returns to schooling in Indonesia, using a new school construction program across regions as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. Hereafter, we refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
@@
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several recent papers impose this assumption with settings of a non-staggered treatment; see, e.g., de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021). [^1] When treatment is binary and there exist only two periods, one can explicitly introduce the treatment path in potential outcomes because it takes only four values. We therefore consider this case in Appendix B as an extension. One can also weaken Assumption 1 in other cases such as non-binary, ordered treatments or multiple time periods, but this would requires the cumbersome notation and complicates the definition of our target parameter, thus is beyond scope of this paper.
Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$, which describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$. Specifically, the first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group. Following to the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ to be the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ to be the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ to be the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ to be the defiers.
Henceforth, we keep Assumption 1. In the next section, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
@@ 2.2   Target parameter in two time periods
Our target parameter in $2\times 2$ DID-IV designs is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
@@
This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$. In the DID-IV literature, de Chaisemartin (2010) consider the same target parameter, and Hudson et al. (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$. The LATET has been also proposed in heterogeneous effects IV models with binary instrument; see, e.g., Słoczyński et al. (2022) and Słoczyński (2020).
We define the LATET as our target parameter for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)). Duflo (2001), for instance, are interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”. To answer this question, the author exploits a new school construction program across regions as an instrument for education attainment. In this context, the LATET would be a policy-relevant parameter in the presence of heterogeneous treatment effects. Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT). In section 2.5, we show that both causal parameters measure the treatment effects among the units who are affected by instrument (policy shock) and belong to an exposed group $(E_{i}=1)$.
@@
de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting we are considering here222de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. We therefore rewrite their target parameter by the notation we use in this article..
@@
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). As we will show in Appendix E, this parameter is different from the LATET and can be decomposed into a weighted average of two different causal parameters. In section 2.6, we provide a detailed discussion of the connections between this article and de Chaisemartin and D’Haultfœuille (2018).
@@ 2.3   Identification assumptions in two time periods
This section establishes DID-IV designs in a simple setting with two groups and two periods. In this setting, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
@@
Following to the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand. This estimand is obtained from the IV regression of the outcome on the treatment where the group and post-time dummies are included instruments and the interaction of the two is excluded instrument.
@@
This assumption requires that the instrument path does not directly affect potential outcomes other than through treatment. This assumption is common in the IV literature; see e.g., Imbens and Angrist (1994) and Abadie (2003). In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) impose the similar assumption without introducing the instrument path.
@@
Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path were $z$:
@@
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes. Note that we can not observe different potential outcomes for the same unit at the same time. For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument over two time periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
@@
This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way. This assumption implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) make the same assumption. Hereafter, we consider the type of monotonicity assumption which rules out the existence of the defiers $DF^{Z}$.
@@
Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group. This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex-ante not known for all the units in an exposed group. This assumption is ”hidden assumption” in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al. (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways. Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume the no anticipation for all units $i$. Roth et al. (2023) take the intermediate approach: they assume the no anticipation for the treated units. Our no anticipation assumption on potential treatment choices is in line with that of Roth et al. (2023).
Next, we impose the parallel trends assumptions in the treatment and the outcome. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) assume the similar assumptions.
@@
Assumption 6 is a parallel trends assumption in the outcome. This assumption requires that the evolution of the unexposed outcome is the same on average between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of the policy shock.
@@
The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the local average treatment effects on the treated (LATET) in period $1$.
@@
If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$, that is,
@@
We now provide the intuition behind $2\times 2$ DID-IV designs. As we already noted, the Wald-DID estimand scales the DID estimand of the outcome by the DID estimand of the treatment. Motivated by this observation, in $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1). First, in the denominator, we use the DID identification strategy in order to identify the direct effect of the instrument on the treatment: we recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment). Next, in the numerator, we use the DID identification strategy in order to identify the indirect effect of the instrument on the outcome through treatment: we recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome). Finally, by scaling the indirect effect obtained from the numerator by the direct effect obtained from the denominator, we can capture the LATET in period $t=1$.
Figure 1: DID-IV identification strategy in two periods and two groups. Notes: This figure plots the evolution of the expectation of the treatment (First stage DID) and the outcome (Reduced form DID) for exposed and unexposed groups respectively. The dotted line represents the counterfactual trends of the expectation of the treatment and the outcome in an exposed group. The effects of the instrument on the treatment and the outcome are $0.2$ and $2$ respectively, and the local average treatment effect on the treated is $10$.

### v4

@@ 1.1   Related literature
Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018)) and contributes to this literature in three important ways.
@@
First, this paper investigates the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption [^1] In biometrics, Ye et al. (2023) also consider the ”instrumented difference-in-differences” design. However, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to identify this parameter.. Hudson et al. (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings. Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs. The settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
@@
In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al. (2017). Specifically, we introduce the path of the instrument in $2\times 2$ DID-IV designs and uncover the ”hidden assumption” in the previous literature. This assumption requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument. We also investigate the selection mechanism behind $2\times 2$ DID-IV designs, and clarify the interpretation of the parallel trends assumption in the outcome.
@@
Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs. Specifically, we show that the identifying assumptions in Fuzzy DID designs ex ante requires strong restrictions on treatment adoption behavior across units, and the target parameter in Fuzzy DID designs can be decomposed into a weighted average of two different causal parameters. The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicates the relative advantages of DID-IV over Fuzzy DID.
@@
Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units. In practice, to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings. Our staggered DID-IV designs ensure the theoretical validity of the underlying identification strategy behind these studies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
@@
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g. Black et al. (2005), Lundborg et al. (2014)). In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time. The proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
@@
Our paper is also related to a growing DID literature. Several recent studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al. (2021); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)). All of them have documented that regression coefficients by conventional approaches may fail to properly summarize treatment effects under DID designs. Some of them have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity. None of them, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in policy adoption timing across units as an instrument for treatment. This paper formalizes the rationale behind using the TWFEIV regression as an instrumented difference-in-differences design, and provide a reliable estimation method in multiple periods settings. Miyaji (2024) builds on this work to study the properties of the TWFEIV estimand under staggered DID-IV design and points out the issue of implementing this design via TWFEIV regressions.
@@ 2   DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings. We first establish the target parameter and the identifying assumptions in this design. We then uncover the selection mechanism behind this design, and clarify the interpretation of the parallel assumption in the outcome. We next describe the relationship between DID and DID-IV, indicating that DID-IV is DID with the noncompliance of the assigned treatment. At the end of this section, we also describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
@@ 2.1   Set up
We introduce the notation we use throughout Section 2. We consider a panel data setting with two periods and $N$ units. For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support. For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receive the treatment. Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$. Throughout Section 2, we assume that $\{Y_{i,0},Y_{i,1},D_{i,0},D_{i,1},Z_{i,0},Z_{i,1}\}_{i=1}^{N}$ are independent and identically distributed (i.i.d).
We introduce the path of the treatment and the instrument. Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ the instrument path. We assume no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$; we refer to this as a sharp assignment of the instrument. We denote $E_{i}\in\{0,1\}$ as the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group). In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment. For instance, Duflo (2001) estimates returns to schooling in Indonesia, using a new school construction program across regions as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
@@
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several studies have used this assumption in non-staggered treatment settings (e.g. de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)). [^2] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes because it takes only four values. We therefore consider this case in Appendix B as an extension. One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatments, or multiple time periods, but it would require cumbersome notations and complicate the definition of the target parameter, going beyond scope of this paper.
Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$, which describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$. The first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group. Following the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ as the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ as the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ as the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ as the defiers.
Henceforth, we keep Assumption 1. In the next subsection, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
@@ 2.2   Target parameter in two time periods
In $2\times 2$ DID-IV designs, our target parameter is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
@@
This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$. In the DID-IV literature, de Chaisemartin (2010) consider the same target parameter, and Hudson et al. (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$. The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g. Słoczyński (2020); Słoczyński et al. (2022)).
We define LATET as the target parameter for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)). Duflo (2001), for instance, is interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”. To answer this question, the author exploits a new school construction program across regions as an instrument for education attainment. In this context, the LATET would be a policy-relevant parameter in the presence of heterogeneous treatment effects. Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT). Subsection 2.5 shows that both causal parameters measure the treatment effects among the units affected by instrument (policy shock) and belonging to an exposed group $(E_{i}=1)$.
@@
de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting as here333de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. Therefore, we rewrite their target parameter by the notation used in this article..
@@
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). As shown in Appendix E, this parameter turns out to be different from the LATET and can be decomposed into a weighted average of two different causal parameters. In section 2.6, we provide a detailed discussion of the connections between this article and de Chaisemartin and D’Haultfœuille (2018).
@@ 2.3   Identification assumptions in two time periods
This subsection establishes DID-IV designs in a simple setting with two groups and two periods. In this setting, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
@@
Following the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand. This estimand is obtained from the IV regression of the outcome on the treatment where the group and post-time dummies are included instruments and the interaction of the two is excluded instrument.
@@
This assumption requires that the instrument path does not directly affect potential outcomes other than through treatment. This assumption is common in the IV literature; see e.g., Imbens and Angrist (1994) and Abadie (2003). In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) impose a similar assumption without introducing the instrument path.
@@
Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path is $z$:
@@
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes. Note that we can not observe different potential outcomes for the same unit at the same time. For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument during the two periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
@@
This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way. It implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) make the same assumption. Hereafter, we consider the type of monotonicity assumption that rules out the existence of the defiers $DF^{Z}$.
@@
Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group. This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex-ante not known for all the units in an exposed group. This is ”hidden assumption” in the DID-IV literature: de Chaisemartin (2010) and Hudson et al. (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways. Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume it for all units $i$. Roth et al. (2023) take the intermediate approach: they adopt the no anticipation assumption for the treated units. In this paper, the no anticipation assumption on potential treatment choices is in line with that of Roth et al. (2023).
Next, we impose the parallel trends assumptions in the treatment and the outcome. de Chaisemartin (2010) and Hudson et al. (2017) make the similar assumptions.
@@
Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
@@
The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the LATET in period $1$.
@@
If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
@@
We now provide the intuition behind $2\times 2$ DID-IV designs. As we already noted, the Wald-DID estimand scales the DID estimand of the outcome by the DID estimand of the treatment. Motivated by this observation, under $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1). First, in the denominator of the Wald-DID estimand, we use the DID identification strategy to identify the direct effect of the instrument on the treatment. We recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment). Next, in the numerator of the Wald-DID estimand, we use the DID identification strategy to identify the indirect effect of the instrument on the outcome through treatment. We recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome). Finally, by scaling the indirect effect obtained from the numerator by the direct effect obtained from the denominator, we can capture the LATET in period $t=1$.
Figure 1: DID-IV identification strategy in two periods and two groups. Notes: This figure plots the evolution of the expectation of the treatment (First-stage DID) and the outcome (Reduced-form DID) for exposed and unexposed groups, respectively. The dotted line represents the counterfactual trends of the expectation of the treatment and the outcome in an exposed group. The effects of the instrument on the treatment and the outcome are $0.2$ and $2$ respectively, and the local average treatment effect on the treated is $10$.